# Accurate close interactions of Stokes spheres using lubrication-adapted image systems

### v1

@@ Abstract
Stokes flows with near-touching rigid particles induce near-singular lubrication forces under relative motion, making their accurate numerical treatment challenging. With the aim of controlling the accuracy with a computationally cheap method, we present a new technique that combines the method of fundamental solutions (MFS) with the method of images. For rigid spheres, we propose to represent the flow using Stokeslet proxy sources on interior spheres, augmented by lines of image sources adapted to each near-contact to resolve lubrication. Source strengths are found by a least-squares solve at contact-adapted boundary collocation nodes. We include extensive numerical tests, and validate against reference solutions from a well-resolved boundary integral formulation. With less than 60 image sources per particle per contact, we show controlled uniform accuracy to three relative digits in surface velocities, and up to five digits in particle forces and torques, for all separations down to a thousandth of the radius. In the special case of flows around fixed particles, the proxy sphere alone gives controlled accuracy. A one-body preconditioning strategy allows acceleration with the fast multipole method, hence close to linear scaling in the number of particles. This is demonstrated by solving problems of up to 2000 spheres on a workstation using only 700 proxy sources per particle.
@@ 1 Introduction
The approach of this paper is, broadly speaking, a method of fundamental solutions (MFS). The solution flow $\boldsymbol{u}$ is written as a linear combination of fundamental solutions to the Stokes PDE, located in the non-physical region (the sphere interiors), whose coefficients are found by enforcing the boundary data at surface collocation points in the least-squares sense. This idea has a long history, and sometimes also goes under the names of the charge simulation method, the method of auxiliary sources, or proxy-sources [42, 43, 44, 45, 46, 47, 48]. A choice of locations and collocation points leading to small errors depends on the domain and data, is heuristic, and is not always possible to determine. However, there are rigorous convergence results in certain smooth geometries, such as the disk with source locations on a concentric circle [49], and when the solution may be analytically continued as a PDE solution up to the source curve [50]. The MFS has the disadvantage of ill-conditioned rectangular linear systems, when compared against the well-conditioned square ones arising from discretizing 2nd-kind Fredholm BIE. However, in special geometries such as close-to-touching spheres, we will see that the advantage of much smaller numbers of unknowns at a given accuracy makes the MFS rather favorable.
@@ 2.1 The basic discretization for one particle
Figure 2: Panel (a) displays the basic discretization, with the original source points $\boldsymbol{y}_{j}$, $j=1,\dots,N$, on the radius $R_{p}$ proxy-surface in blue, and collocation points $\boldsymbol{x}_{i}$, $i=1,\dots,M$, in black on the unit sphere surface. Both source and collocation points are obtained from spherical design nodes. Panels (b) and (c) display the image-enhanced grid of source and collocation points for a pair of close-to-touching spheres. Additional image points are distributed along a line segment in the interior of each sphere, while additional collocation points are distributed on two spherical caps surrounding the point of closest approach to another particle “above” the line of images, as in panel (b). If a particle is closely interacting with more than one neighbor, one region of extra source and collocation points of this type is added for every near-contact. In panel (c), collocation points are indicated in black and image points in red, generated via (26)-(27).
@@ 2.2 One-body preconditioned iterative solution for many particles
$\hat{\boldsymbol{G}}{\boldsymbol{\mu}}:=\begin{bmatrix}\boldsymbol{I}&\boldsymbol{G}^{(12)}\boldsymbol{B}^{+}&\dots&\boldsymbol{G}^{(1P)}\boldsymbol{B}^{+}\\ \boldsymbol{G}^{(21)}\boldsymbol{B}^{+}&\boldsymbol{I}&\ldots&\ldots\\ \vdots&\vdots&\ddots&\vdots\\ \boldsymbol{G}^{(P1)}\boldsymbol{B}^{+}&\ldots&\ldots&\boldsymbol{I}\end{bmatrix}\begin{bmatrix}\boldsymbol{\mu}^{(1)}\\ \boldsymbol{\mu}^{(2)}\\ \vdots\\ \boldsymbol{\mu}^{(P)}\end{bmatrix}=\begin{bmatrix}\boldsymbol{u}^{(1)}\\ \boldsymbol{u}^{(2)}\\ \vdots\\ \boldsymbol{u}^{(P)}\end{bmatrix}.$ (9)
@@
Upon GMRES convergence, the vector $\boldsymbol{\mu}$ is again via (11) transformed back to the source strength vector $\boldsymbol{\lambda}$. The latter can then be used for solution velocity or pressure field evaluation throughout the fluid.
@@ 2.3 Image enhancement for problems with strong lubrication forces
$\mathbb{T}_{ijk}(\boldsymbol{r})=-6\frac{r_{i}r_{j}r_{k}}{\|\boldsymbol{r}\|^{5}}$ (14)
@@
$\mathbb{R}(\boldsymbol{r})=\frac{1}{\|\boldsymbol{r}\|^{3}}\begin{bmatrix}0&r_{3}&-r_{2}\\ -r_{3}&0&r_{1}\\ r_{2}&-r_{1}&0\end{bmatrix}.$ (16)
@@
$\boldsymbol{u}(\boldsymbol{x})=\frac{1}{8\pi\mu}\mathbb{R}(\boldsymbol{x}-\boldsymbol{y})\boldsymbol{t},$ (17)
@@
$\boldsymbol{u}(\boldsymbol{x})=\frac{1}{4\pi}\mathbb{D}(\boldsymbol{x}-\boldsymbol{y})\boldsymbol{d},$ (18)
@@
$\mathbb{D}(\boldsymbol{r})=-\frac{\boldsymbol{I}}{\|\boldsymbol{r}\|^{3}}+3\frac{\boldsymbol{r}\boldsymbol{r}^{T}}{\|\boldsymbol{r}\|^{5}}.$ (19)
@@ 3 Parameters for the basic concentric proxy sphere
This section serves to optimize the MFS proxy sphere parameters (the number of proxy-sources $N$ and the radius of the proxy-surface $R_{p}$) used in later experiments. No image points are used, and the number of collocation points is fixed at $M\approx 1.2N$, a ratio found by a convergence study in $M$ at fixed $N$.
@@ 5.2 Validating the pairwise-adaptive strategy in clusters of spheres
Figure 14: Comparison to results from a BIE formulation for the tetrahedron test geometry of Figure 13(a); forces and torques determined with the image enhanced grid are compared to different sets of reference data. For small $\delta$, the BIE-reference is inaccurate due to the under-resolved peak in the layer density. This causes increasing errors for decreasing $\delta$ in comparisons of coarse MFS results both to solutions determined with coarse and fine BIE-grids (see details in main text), but with different onsets of the increase. For larger $\delta$, however, the error level in the forces and torques from the self-convergence test is similar in magnitude to the error in the comparison to a solution computed with BIE. For BIE, the QBX-tolerance is set to $10^{-6}$, and hence we do not expect smaller errors to be detected. Red and blue represent the same type of comparisons, but using different reference solutions: for each separation distance $\delta$, the experiment is repeated 20 times with randomly generated boundary data, and the min, max and mean of $\epsilon_{\text{FT}}$ are visualized. Test problems generated with a unit force or torque are labeled ’bc A’ and ’bc B’. The self-convergence data is the same as presented in Figure 13(b).
@@ 6 Conclusions and outlook
A natural extension to the work reported here is to the mobility problem. We will report on a well-conditioned MFS formulation for the mobility problem separately. A different future direction is to apply a similar framework to polydisperse spheres, and to non-spherical smooth particles. For simple particles with separations where proxy surfaces suffice to resolve interactions, such a generalization is straightforward. However, image representations are harder to generalize and form a topic of future research. In Section 5.3, we note that the number of GMRES iterations grows with decreasing particle gaps; future work includes investigating suitable preconditioning. As for the size of the systems considered in this work, all tests fit on a standard workstation. Larger systems could be studied via a distributed memory implementation and FMM, as in [30].
@@ Acknowledgements
Broms and Tornberg acknowledge the support from the Swedish Research Council: grant no. 2019-05206 and the research environment grant INTERFACE (biomaterials), no. 2016-06119. The Flatiron Institute is a division of the Simons Foundation. Broms appreciates the research visit at CCM Flatiron, which served as a catalyst for the initial ideas in this work.

### v2

@@ Abstract
Stokes flows with near-touching rigid particles induce near-singular lubrication forces under relative motion, making their accurate numerical treatment challenging. With the aim of controlling the accuracy with a computationally cheap method, we present a new technique that combines the method of fundamental solutions (MFS) with the method of images. For rigid spheres, we propose to represent the flow using Stokeslet proxy sources on interior spheres, augmented by lines of image sources adapted to each near-contact to resolve lubrication. Source strengths are found by a least-squares solve at contact-adapted boundary collocation nodes. We include extensive numerical tests, and validate against reference solutions from a well-resolved boundary integral formulation. With less than 60 additional image sources per particle per contact, we show controlled uniform accuracy to three relative digits in surface velocities, and up to five digits in particle forces and torques, for all separations down to a thousandth of the radius. In the special case of flows around fixed particles, the proxy sphere alone gives controlled accuracy. A one-body preconditioning strategy allows acceleration with the fast multipole method, hence close to linear scaling in the number of particles. This is demonstrated by solving problems of up to 2000 spheres on a workstation using only 700 proxy sources per particle.
@@ 1 Introduction
The approach of this paper is, broadly speaking, a method of fundamental solutions (MFS). The solution flow $\boldsymbol{u}$ is written as a linear combination of fundamental solutions to the Stokes PDE, located in the non-physical region (the sphere interiors), whose coefficients are found by enforcing the boundary data in the least-squares sense at a similar number of surface collocation points. This idea has a long history, and sometimes also goes under the names of the charge simulation method, the method of auxiliary sources, or proxy-sources [42, 43, 44, 45, 46, 47, 48]. A choice of locations and collocation points leading to small errors depends on the domain and data, is heuristic, and is not always possible to determine. However, there are rigorous convergence results in certain smooth geometries, such as the disk with source locations on a concentric circle [49], and when the solution may be analytically continued as a PDE solution up to the source curve [50]. The MFS has the disadvantage of involving ill-conditioned rectangular linear systems, when compared against the well-conditioned square ones arising from discretizing 2nd-kind Fredholm BIEs. In special geometries such as close-to-touching spheres, however, we will see that the advantage of much smaller numbers of unknowns at a given accuracy makes the MFS rather favorable compared to BIEs.
@@ 1.1 Contribution and outline
Remark 1 (Recent Stokes MFS work).
Since the preprint of this paper appeared on arXiv, another MFS Stokes paper appeared by Jordan & Lockerby [58]. They present results for the resistance and mobility problems, including nonspherical particles. However, in contrast to our work, only problems with weak lubrication effects (large separations) are treated. They apply Gauss–Seidel iterations to account for particle-particle effects, giving a $\mathcal{O}(P^{2})$ cost, contrasting the linear scaling obtained via the FMM in our work. Results in [58] are shown for particles along a line, with interactions greatly simplified beyond a certain distance. In contrast, we present a method with full interactions and test it with space-filling configurations.
@@ 2.1 The basic discretization for one particle
The ratio $M/N\approx 1.2$ was found by a convergence study, and is typical for MFS applications [44, 50]. We do not know of rigorous analysis of this ratio, apart from the case $M/N=1$ for the 2D Laplace equation [49]. Intuitively, a ratio greater than $1$ is beneficial because the potential generated by somewhat nearby proxy points contains frequency components higher than merely the Nyquist frequency associated with their spacing.
@@
Figure 2: Panel (a) displays the basic discretization, with the original source points $\boldsymbol{y}_{j}$, $j=1,\dots,N$, on the radius $R_{p}$ proxy-surface in blue, and collocation points $\boldsymbol{x}_{i}$, $i=1,\dots,M$, in black on the unit sphere surface. Both source and collocation points are obtained from spherical design nodes. Panels (b) and (c) display the image-enhanced grid of source and collocation points for a pair of close-to-touching spheres. Additional image points are distributed along a line segment in the interior of each sphere, while additional collocation points are distributed “above” each line of images. The collocation points are located on two spherical caps determined by angles $\beta_{1}$ and $\beta_{2}$ surrounding the point of closest approach to another particle, as in panel (b). If a particle is closely interacting with more than one neighbor, one region of extra source and collocation points of this type is added for every near-contact. In panel (c), collocation points are indicated in black and image points in red, generated via (26)-(27).
@@ 2.2 One-body preconditioned iterative solution for many particles
$\hat{\boldsymbol{G}}{\boldsymbol{\mu}}:=\begin{bmatrix}\boldsymbol{I}&\boldsymbol{G}^{(12)}\boldsymbol{B}^{+}&\dots&\boldsymbol{G}^{(1P)}\boldsymbol{B}^{+}\\ \boldsymbol{G}^{(21)}\boldsymbol{B}^{+}&\boldsymbol{I}&\ldots&\ldots\\ \vdots&\vdots&\ddots&\vdots\\ \boldsymbol{G}^{(P1)}\boldsymbol{B}^{+}&\ldots&\ldots&\boldsymbol{I}\end{bmatrix}\begin{bmatrix}\boldsymbol{\mu}^{(1)}\\ \boldsymbol{\mu}^{(2)}\\ \vdots\\ \boldsymbol{\mu}^{(P)}\end{bmatrix}={\begin{bmatrix}\boldsymbol{b}^{(1)}\\ \boldsymbol{b}^{(2)}\\ \vdots\\ \boldsymbol{b}^{(P)}\end{bmatrix}.}$ (9)
@@
If direct summation is used, of course steps 2 and 3 could be combined, but in order to tackle the largest problems we will focus on the use of the FMM where the steps must be separate. Upon GMRES convergence, the vector $\boldsymbol{\mu}$ is again via (11) transformed back to the source strength vector $\boldsymbol{\lambda}$. The latter can then be used for solution velocity or pressure field evaluation throughout the fluid.
@@ 2.3 Image enhancement for problems with strong lubrication forces
$\mathbb{T}_{ijk}(\boldsymbol{r})=-\frac{1}{8\pi\mu}\frac{3r_{i}r_{j}r_{k}}{\|\boldsymbol{r}\|^{5}}$ (14)
@@
$\mathbb{R}(\boldsymbol{r})=\frac{1}{8\pi\mu}\frac{1}{\|\boldsymbol{r}\|^{3}}\begin{bmatrix}0&r_{3}&-r_{2}\\ -r_{3}&0&r_{1}\\ r_{2}&-r_{1}&0\end{bmatrix}.$ (16)
@@
$\boldsymbol{u}(\boldsymbol{x})=\mathbb{R}(\boldsymbol{x}-\boldsymbol{y})\boldsymbol{t},$ (17)
@@
$\boldsymbol{u}(\boldsymbol{x})=\mathbb{D}(\boldsymbol{x}-\boldsymbol{y})\boldsymbol{d},$ (18)
@@
$\mathbb{D}(\boldsymbol{r})=\frac{1}{4\pi}\left(-\frac{\boldsymbol{I}}{\|\boldsymbol{r}\|^{3}}+3\frac{\boldsymbol{r}\boldsymbol{r}^{T}}{\|\boldsymbol{r}\|^{5}}\right).$ (19)
@@ 3 Parameters for the basic concentric proxy sphere
This section serves to optimize the MFS proxy sphere parameters (the number of proxy-sources $N$ and the radius of the proxy-surface $R_{p}$) used in later experiments. No image points are used, and the number of collocation points is fixed at $M\approx 1.2N$, as discussed in Section 2.
@@ 5.2 Validating the pairwise-adaptive strategy in clusters of spheres
Figure 14: Comparison to results from a BIE formulation for the tetrahedron test geometry of Figure 13(a); forces and torques determined with the image enhanced grid are compared to different sets of reference data. For small $\delta$, the BIE-reference is inaccurate due to the under-resolved peak in the layer density. This causes increasing errors for decreasing $\delta$ in comparisons of coarse MFS results both to solutions determined with coarse and fine BIE-grids (see details in main text), but with different onsets of the increase. For larger $\delta$, however, the error level in the forces and torques from the self-convergence test is similar in magnitude to the error in the comparison to a solution computed with BIE. For BIE, the QBX-tolerance is set to $10^{-6}$, which together with the GMRES tolerance restrict how small errors that can be detected for larger $\delta$. Red and blue represent the same type of comparisons, but using different reference solutions: for each separation distance $\delta$, the experiment is repeated 20 times with randomly generated boundary data, and the min, max and mean of $\epsilon_{\text{FT}}$ are visualized. Test problems generated with a unit force or torque are labeled ’bc A’ and ’bc B’. The self-convergence data is the same as presented in Figure 13(b).
@@ 6 Conclusions and outlook
A natural extension to the work reported here is to the mobility problem. We report on a well-conditioned MFS formulation for the mobility problem separately in [75]. In that work, we also extend to non-spherical smooth particles in the setting where proxy surfaces suffice to resolve interactions. It is a topic of future research whether image representations could be useful for non-spherical particles. In Section 5.3, we note that the number of GMRES iterations grows with decreasing particle gaps. Future work includes investigating suitable preconditioning or the use of a fast direct solver. As for the size of the systems considered in this work, all tests fit on a standard workstation. Larger systems could be studied via a distributed memory implementation and FMM, as in [30].
@@ Acknowledgements
Broms and Tornberg acknowledge the support from the Swedish Research Council: grant no. 2019-05206 and the research environment grant INTERFACE (biomaterials), no. 2016-06119. The Flatiron Institute is a division of the Simons Foundation. Broms appreciates the research visit at CCM Flatiron, which served as a catalyst for the initial ideas in this work. We would also like to thank the referees for their thorough reading of the manuscript and for their valuable comments which have helped us to improve the text.